OpenStudy (anonymous):

How would I write an equation of a line in slope-intercept form with a slope of -4 and passing through the point (0, -3)? this is stumping me?

OpenStudy (texaschic101):

slope intercept form : y = mx + b......m is your slope and b is your y intercept y = mx + b slope(m) = -4 y intercept(b) = -3 so just sub y = -4x + (-3)....or y = -4x - 3

OpenStudy (anonymous):

awesome thank you very much! that made it easy to understand!

OpenStudy (texaschic101):

good to hear :)